# Massive Supervised Fine-tuning Experiments Reveal How Data, Layer, and Training Factors Shape LLM Alignment Quality

### v1

@@ Abstract
Supervised fine-tuning (SFT) is a critical step in aligning large language models (LLMs) with human instructions and values, yet many aspects of SFT remain poorly understood. We trained a wide range of base models on a variety of datasets including code generation, mathematical reasoning, and general-domain tasks, resulting in 1,000+ SFT models under controlled conditions. We then identified the dataset properties that matter most and examined the layer-wise modifications introduced by SFT. Our findings reveal that some training–task synergies persist across all models while others vary substantially, emphasizing the importance of model-specific strategies. Moreover, we demonstrate that perplexity consistently predicts SFT effectiveness—often surpassing superficial similarity between trained data and benchmark—and that mid-layer weight changes correlate most strongly with performance gains. We will release these 1,000+ SFT models and benchmark results to accelerate further research.
@@
Yuto Harada1,2*thanks: Equal Contribution., Yusuke Yamauchi1,2*11footnotemark: 1, Yusuke Oda1,3, Yohei Oseki2, Yusuke Miyao1,2thanks: Corresponding authors: Yusuke Miyao and Yu Takagi, Yu Takagi422footnotemark: 2 1NII LLMC, 2The University of Tokyo, 3NAIST, 4Nagoya Institute of Technology, {harada-yuto, yamauchi_y}@nii.ac.jp yusuke@is.s.u-tokyo.ac.jptakagi.yu@nitech.ac.jp
@@ 1 Introduction
Recent advances in large language models (LLMs) have greatly improved natural language understanding and generation. However, purely pre-trained LLMs often fail to align with human intentions or specific tasks (Ouyang et al., 2022), prompting increasing focus on alignment techniques. Supervised fine-tuning (SFT) trains models to follow human instructions, and remains widely used and effective for improving downstream performance (Wei et al., ; Guan et al., 2024).
Although recent works have explored how model size and training-data characteristics influence downstream tasks in the context of SFT (Jin and Ren, 2024; Dong et al., 2024), large-scale research specifically examining which aspects of SFT datasets benefit different base models remains limited. While some studies compare or analyze publicly available models (Oyama et al., 2025), these are not controlled experiments and often introduce biases—such as favoring certain model families. Consequently, it remains unclear how SFT of various models on different datasets affects benchmark performance, how relationships among datasets and benchmarks vary across models, and which internal weights are most responsible for these effects. Furthermore, there are several SFT training approaches including Low-Rank Adaptation (LoRA) (Hu et al., 2022), and there is ongoing debate about the optimal amount of data required (Zhou et al., 2024; Chen et al., 2023); however, there has yet to be a comprehensive, quantitative comparison. Hence, a comprehensive examination of these issues on SFT is urgently needed.
@@
In this study, we trained twelve diverse base models on multiple datasets spanning different domains, creating a large suite of SFT models that we subsequently evaluated on a broad range of tasks (Figure 1). Specifically, we address the following Research Questions (RQs):
@@
    How do models, training data, and benchmarks interact with one another? Do certain training datasets consistently enhance benchmark performance across a variety of models, or does each model exhibit its own distinct preferences? Likewise, do relationships among different datasets and benchmarks remain the same across models?
@@
    Which layers in the model are most critical for SFT—are there universal patterns across different models?
@@
    How do various factors debated in SFT—such as different training methods, sample sizes, and cross-lingual transfer—impact performance?
The main contributions of this work can be summarized as follows:
@@
We find that training data with lower perplexity for the base model consistently leads to greater improvements in downstream performance. In contrast, factors once considered crucial—such as content similarity between training and evaluation data or tokenizer compatibility—do not exhibit as strong an effect as perplexity.
@@
Projecting the log-likelihood vectors of fine-tuned models into a common latent space lets us compare diverse training dynamics in one coordinate system. The resulting map shows that the global layout is determined by model family rather than training corpus, that checkpoints from successive epochs converge toward a shared instruction-following region, that enlarging the instruction set from 1k to 20k nudges models only slightly outward from this centre, and that LoRA trajectories almost perfectly overlap those of full-parameter tuning.
@@
All fine-tuned models produced in this study will be publicly released. We expect this comprehensive set of models serves to accelerate deeper investigations of SFT and to foster rapid progress in the field.
@@ 2 Related Work
The role of training data characteristics in SFT has been highlighted in many prior studies. For instance, mixing code-generation data has been suggested to enhance a model’s reasoning and logical abilities (Dong et al., 2024). Similarly, incorporating instruction data that includes procedural knowledge could improve mathematical reasoning (Ruis et al., 2024). Furthermore, considering task relevance when selecting datasets can lead to more robust general performance (Huang et al., 2024; Zhang et al., 2024).
While early work focused on how to fine-tune—comparing full-parameter updates against LoRA (Ivison et al., 2023; Zhuo et al., 2024; Dettmers et al., 2024; Zhao et al., 2024b; Biderman et al., 2024), or debating sample size (Zhou et al., 2024; Zhao et al., 2024a; Chen et al., 2023)—more recent studies have shifted attention to the statistics of the training data itself. For example, Jin and Ren (2024) and Wu et al. (2025) independently show that lower perplexity and moderate sequence length are stronger predictors of SFT success than sheer volume.
@@ 3.2 Training Datasets
We utilized 10 distinct datasets categorized into 4 major groups. Although our base models cover English, Chinese, and Japanese, all training datasets used for SFT are exclusively in English. Specifically, we selected General Tasks: Alpaca(Taori et al., 2023), LIMA(Zhou et al., 2024), and UltraChat(Ding et al., 2023); Coding Tasks: CodeAlpaca(Chaudhary, 2023) and Magicoder(Wei et al., 2024); Math Tasks: OpenMathInstruct(Toshniwal et al., 2024) and MathInstruct(Yue et al., 2023); and Classic NLP Tasks: FLAN(Wei et al., ). The FLAN dataset(Wei et al., ) further consists of 3 subcategories. FLAN Knowledge includes BoolQ(Clark et al., 2019), NaturalQuestions(Kwiatkowski et al., 2019b), and TriviaQA(Joshi et al., 2017). FLAN Reasoning includes ARC-Easy & Challenge(Clark et al., 2018), HellaSwag(Zellers et al., 2019), WinoGrande(Sakaguchi et al., 2019), and PIQA(Bisk et al., 2020). FLAN Comprehension includes QuAC(Choi et al., 2018) and SQuAD v2(Rajpurkar et al., 2018). The categorization of FLAN follows the criteria defined in Dubey et al. (2024); Contributors (2023).
@@ 3.3 Training Settings
Figure 2: a Average of the performance change for diverse benchmarks from the each baseline model after SFT on each training dataset. Each column is min-max scaled to the [$-1,1$] range. b The performance changes visualized for each model individually. c Pairwise correlation matrix of performance changes across all SFT models, with the corresponding hierarchical-clustering dendrogram superimposed. d The cumulative explained variance ratio obtained by applying PCA to all concatenated results from b.
@@ 4 Results
As shown in Figure 4a, there is a clear positive correlation in many tasks and models between lower perplexity (w.r.t. the base model) and improved downstream performance. This implies that data lying in a domain or language distribution already “understood” by the model can be leveraged more effectively in SFT.
Figure 4b reveals modest correlation between the mean token length of a dataset and downstream performance, suggesting that simply using shorter or longer texts does not strongly drive better results. A prior study has reported that longer texts could be important for improved performance (Zhao et al., 2024a), and our findings partially support a straightforward link between text length and outcome quality.
@@
In sum, perplexity relative to the base model emerges as a strong predictor of downstream gains, surpassing factors like token length or broad semantic alignment.
4.3 RQ3. Layer-wise weight changes, their relationship to performance, and the effect of SFT on representational dimensionality.
@@
We then explore how model parameters shift during fine-tuning by analyzing layer-wise weight updates across multiple models. Our goal is to identify which layers are most critical in translating SFT into performance gains.
Figure 5a plots two curves: the blue line is the Pearson correlation between weight-delta magnitude and overall accuracy gain, whereas the orange line shows the raw weight-delta magnitude itself. The orange line grows toward upper layers, yet the blue line peaks in the middle, indicating that the largest edits are not the most consequential ones. Rather, we find that the middle layers exhibit the strongest positive correlation with performance gains.
@@
Figure 5c extends this idea across models: it correlates, for different layer, the weight-change vector of one model with the corresponding vector of every other model. The strongest agreement again lies in the mid-layers, suggesting that SFT enforces a shared instruction-following mechanism across models.
@@
Finally, we consider additional aspects of SFT, including LoRA versus full-parameter tuning, the effect of sample size, and cross-lingual transfer—each potentially influencing the final performance.
@@
When points are coloured by model (Fig. 6a) the clusters group almost perfectly by architecture, whereas colouring by training data produces only weak separation. Thus the inductive biases of the base model outweigh the specific SFT corpus in determining the final representation.
@@
Colouring by training-set size separates models trained on 20k samples from those trained on 1k samples. The 20k-sample–trained points occupy the outer rim of the manifold more often, whereas the 1k-sample–trained points cluster nearer the core. Thus a compact 1k instruction set already supplies sufficient signal for effective instruction-tuning, while scaling up to 20k samples can sometimes pull the representation away from the optimum. Indeed, our quantitative evaluations showed no consistent accuracy advantage for the 20k-sample models over their 1k-sample counterparts.
@@ 5 Discussion and Conclusion
We conducted a comprehensive set of SFT experiments involving multiple 7B-scale base models, diverse training datasets, and a wide array of downstream tasks. Our analysis revealed that, while certain dataset–task synergies are observed consistently across models, their effects can vary greatly depending on the specific model in question. Notably, perplexity emerged as a particularly robust predictor of SFT success, outperforming both topic similarity and average sequence length. Furthermore, mid-layer weight changes were found to correlate most strongly with performance improvements, indicating that critical adaptations often take place in these layers. By embedding every model checkpoint into a common latent space, we found that (i) model architecture exerts a stronger influence than the SFT corpus, (ii) training epochs drive diverse runs toward a shared instruction-compatible region, (iii) large instruction sets tend to relocate models toward the periphery—often reducing accuracy relative to leaner sets—and (iv) LoRA trajectories almost coincide with full-parameter ones, diverging only slightly on the periphery; this mirrors the small but systematic trade-off we observed between knowledge-heavy tasks (full-parameter advantaged) and open-ended QA (LoRA advantaged).
@@
Discovering the importance of mid-layer changes could reshape fine-tuning strategies. Updating only the mid-layers, or monitoring their changes more closely, could provide more efficient or interpretable SFT. Furthermore, observing common mid-layer change patterns across models indicates a potentially shared mechanism for task-related knowledge acquisition.
@@
The present study focuses on around 7B-parameter models, leaving open the question of whether similar patterns hold for larger models. We use about 10 popular training datasets for SFT, possibly limiting generality for highly specialized tasks or broader multilingual corpora. While perplexity proved insightful, it can fluctuate based on tokenizer design and base training distributions, indicating a need for more nuanced measures.
@@
We will release over 1,000 fine-tuned models as part of this study. While we do not anticipate major risks, we acknowledge the potential for misuse—such as generating harmful or misleading content. To mitigate this, all released models will include a responsible use clause and detailed model cards describing limitations. We encourage responsible use for academic research only.
@@ Appendix B Description of Training Datasets
Classic NLP FLAN Knowledge Open-Orca/FLAN 226,575 21.69
@@
FLAN Collection (Wei et al., ) is the remix file flan2021_zsnoopt_submix_data.json. Specifically, it corresponds to the FLAN-2021 sub-mix and employs the zero-shot, no-options template variant (i.e., prompts contain only the instruction without in-context examples or candidate options). We follow the taxonomy of Dubey et al. (2024); Contributors (2023) and split the data into three thematic subsets.
@@ Appendix C Preliminary Experiments
This combination of hyperparameters resulted in 96 unique experimental conditions. Each condition was trained for 10 epochs, yielding 960 models per dataset. Given that we utilized six datasets, the total number of trained models reached 5,760.
@@ Appendix D Description of Training Settings
The computational time for fine-tuning on 1k samples varied depending on the model, batch size, and training method, but on average, each run took approximately 30 minutes. To accelerate training, we employed Flash Attention 2 (Dao, 2023) and DeepSpeed (Rasley et al., 2020) for all models.

### v2

@@ Abstract
Supervised fine-tuning (SFT) is a critical step in aligning large language models (LLMs) with human instructions and values, yet many aspects of SFT remain poorly understood. We trained a wide range of base models on a variety of datasets including code generation, mathematical reasoning, and general-domain tasks, resulting in 1,000+ SFT models under controlled conditions. We then identified the dataset properties that matter most and examined the layer-wise modifications introduced by SFT. Our findings reveal that some training–task synergies persist across all models while others vary substantially, emphasizing the importance of model-specific strategies. Moreover, we demonstrate that perplexity consistently predicts SFT effectiveness, often surpassing superficial similarity between the training data and the benchmark, and that mid-layer weight changes correlate most strongly with performance gains. We release these 1,000+ SFT models and benchmark results to accelerate further research. All resources are available at https://github.com/llm-jp/massive-sft.
@@
Yuto Harada1,2*thanks: Equal Contribution., Yusuke Yamauchi1,2*11footnotemark: 1, Yusuke Oda1,3, Yohei Oseki1,2, Yusuke Miyao1,2thanks: Corresponding authors: Yusuke Miyao and Yu Takagi, Yu Takagi422footnotemark: 2 1NII LLMC, 2The University of Tokyo, 3NAIST, 4Nagoya Institute of Technology, {harada-yuto, yamauchi_y}@nii.ac.jp yusuke@is.s.u-tokyo.ac.jptakagi.yu@nitech.ac.jp
@@ 1 Introduction
Recent advances in large language models (LLMs) have greatly improved natural language understanding and generation. However, pretrained LLMs often fail to align with human intentions or specific tasks (Ouyang et al., 2022), motivating alignment methods. Supervised fine-tuning (SFT) trains models to follow human instructions and remains a widely used and effective approach for improving downstream performance (Wei et al., 2022; Guan et al., 2024).
Although recent works have explored how model size and training-data characteristics influence downstream tasks in the context of SFT (Jin and Ren, 2024; Dong et al., 2024), large-scale research specifically examining which aspects of SFT datasets benefit different base models remains limited. While some studies compare or analyze publicly available models (Oyama et al., 2025), these are not controlled experiments and often introduce biases, such as favoring certain model families. Consequently, it remains unclear how SFT of various models on different datasets affects benchmark performance, how relationships among datasets and benchmarks vary across models, and which internal weights are most responsible for these effects. Furthermore, there are several SFT training approaches including Low-Rank Adaptation (LoRA) (Hu et al., 2022), and there is ongoing debate about the optimal amount of data required (Zhou et al., 2024; Chen et al., 2023); however, there has yet to be a comprehensive, quantitative comparison. Hence, a comprehensive examination of these issues on SFT is urgently needed.
@@
In this study, we trained twelve base models on multi-domain datasets, produced a large suite of SFT models, and evaluated them across diverse tasks (Figure 1). Specifically, we address the following Research Questions (RQs):
@@
    How do models, training data, and benchmarks interact for downstream performance? Do any training datasets yield consistent gains across models, or are improvements model-specific? Are relationships between datasets and benchmarks stable across models?
@@
    Which layers in the model are most critical for SFT? Are there universal patterns across different models?
@@
    How do factors debated in SFT, including training method, sample size, and cross-lingual transfer, relate to performance?
In summary, our contributions are as follows:
@@
We find that training data with lower perplexity for the base model consistently leads to greater improvements in downstream performance. In contrast, factors once considered crucial, such as content similarity between training and evaluation data or tokenizer compatibility, do not exhibit as strong an effect as perplexity.
@@
Projecting the log-likelihood vectors of fine-tuned models into a common latent space lets us compare diverse training dynamics in one coordinate system. The resulting map shows that the global layout is determined by model family rather than training corpus, that checkpoints from successive epochs converge toward a shared instruction-following region, that enlarging the instruction set from 1k to 20k nudges models only slightly outward from this center, and that LoRA trajectories almost perfectly overlap those of full-parameter tuning.
@@
All fine-tuned models produced in this study are publicly released. We expect this comprehensive set of models to accelerate deeper investigations of SFT and to foster rapid progress in the field.
@@ 2 Related Work
The role of training data characteristics in SFT has been highlighted in many prior studies. For instance, mixing code-generation data has been suggested to enhance a model’s reasoning and logical abilities (Dong et al., 2024). Similarly, incorporating instruction data that includes procedural knowledge could improve mathematical reasoning (Ruis et al., 2024). Furthermore, considering task relevance when selecting datasets can lead to more robust performance (Huang et al., 2024; Zhang et al., 2024).
While early work focused on how to fine-tune, comparing full-parameter updates against LoRA (Ivison et al., 2023; Zhuo et al., 2024; Dettmers et al., 2024; Zhao et al., 2024b; Biderman et al., 2024), or debating sample size (Zhou et al., 2024; Zhao et al., 2024a; Chen et al., 2023). More recent studies have shifted attention to the statistics of the training data itself. For example, Jin and Ren (2024) and Wu et al. (2025) independently show that lower perplexity and moderate sequence length are stronger predictors of SFT success than sheer volume.
@@ 3.2 Training Datasets
We utilized 10 distinct datasets categorized into 4 major groups. Although our base models cover English, Chinese, and Japanese, all training datasets used for SFT are exclusively in English. Specifically, we selected General Tasks: Alpaca (Taori et al., 2023), LIMA (Zhou et al., 2024), and UltraChat (Ding et al., 2023); Coding Tasks: CodeAlpaca (Chaudhary, 2023) and Magicoder (Wei et al., 2024); Math Tasks: OpenMathInstruct (Toshniwal et al., 2024) and MathInstruct (Yue et al., 2023); and Classic NLP Tasks: FLAN (Wei et al., 2022). The FLAN dataset further consists of 3 subcategories. FLAN Knowledge includes BoolQ (Clark et al., 2019), NaturalQuestions (Kwiatkowski et al., 2019b), and TriviaQA (Joshi et al., 2017). FLAN Reasoning includes ARC-Easy & Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), and PIQA (Bisk et al., 2020). FLAN Comprehension includes QuAC (Choi et al., 2018) and SQuAD v2 (Rajpurkar et al., 2018). The categorization of FLAN follows the criteria defined in Dubey et al. (2024); Contributors (2023).
@@ 3.3 Training Settings
Figure 2: a Average of the performance change for diverse benchmarks from the each baseline model after SFT on each training dataset. Each column is min-max scaled to the [1,1-1,1] range. b The performance changes visualized for each model individually. c Pairwise correlation matrix of performance changes across all SFT models, with the corresponding hierarchical-clustering dendrogram superimposed. d The cumulative explained variance ratio obtained by applying PCA to all concatenated results from b.
@@ 4 Results
As shown in Figure 4a, there is a clear negative correlation in many tasks and models between lower perplexity (w.r.t. the base model) and improved downstream performance. This implies that data lying in a domain or language distribution already “understood” by the model can be leveraged more effectively in SFT.
Figure 4b reveals a modest correlation between the mean token length of a dataset and downstream performance, suggesting that simply using shorter or longer texts does not strongly drive better results. A prior study has reported that longer texts could be important for performance (Zhao et al., 2024a), and our findings partially support a straightforward link between text length and outcome quality.
@@
In sum, perplexity relative to the base model emerges as a strong predictor of downstream gains, surpassing factors like token length or broad semantic alignment. It thus serves as a practical indicator of model–data compatibility for SFT, even when the base model’s pretraining data are unknown.
4.3 RQ3. Layer-wise weight changes, their relationship to performance, and the effect of SFT on representational dimensionality
@@
We then explore how model parameters shift during fine-tuning by analyzing layer-wise weight updates across multiple models in detail. Our goal is to identify which layers are most critical in translating SFT into performance gains.
Figure 5a plots two curves: specifically, the blue line is the Pearson correlation between weight-delta magnitude and overall accuracy gain, whereas the orange line shows the raw weight-delta magnitude. The orange line grows toward upper layers, yet the blue line peaks in the middle, indicating that the largest edits are not the most consequential ones. Rather, we find that the middle layers in particular exhibit the strongest positive correlation with performance gains.
@@
Figure 5c quantifies cross-model similarity at each layer. We vectorize the SFT weight change per layer for each model, compute its correlation with the corresponding vectors from all other models, and average the resulting pairwise correlations. The strongest agreement again lies in the mid-layers, suggesting that SFT enforces a shared instruction-following mechanism across models.
@@
Finally, we consider additional aspects of SFT, including LoRA versus full-parameter tuning, the effect of sample size, and cross-lingual transfer, each of which may influence final performance.
@@
When points are coloured by model (Fig. 6a) the clusters group almost perfectly by architecture, whereas colouring by training data (Fig. 6b) produces only weak separation. Thus the inductive biases of the base model outweigh the specific SFT corpus in determining the final representation.
@@
Colouring by training‐set size separates models trained on 20k samples from those trained on 1k samples. The 20k-sample–trained points occupy the outer rim of the manifold more often, whereas the 1k-sample–trained points cluster nearer the core. Thus a compact 1k instruction set already supplies sufficient signal for effective instruction-tuning, while scaling up to 20k samples can sometimes pull the representation away from the optimum. Indeed, our quantitative evaluations showed no consistent accuracy advantage for the 20k-sample models over their 1k-sample counterparts.
@@ 5 Discussion and Conclusion
We conducted a comprehensive set of SFT experiments involving multiple 7B-scale base models, diverse training datasets, and a wide array of downstream tasks. Our analysis revealed that, while certain dataset–task synergies are observed consistently across models, their effects can vary greatly depending on the specific model in question. Notably, perplexity emerged as a particularly robust predictor of SFT success, outperforming both topic similarity and average sequence length. Perplexity can be shaped by multiple latent properties of both the data and the model. Accordingly, we view it as a practical proxy for compatibility between the model and the data rather than a causal factor. Identifying the causal drivers behind this association is an important direction for future work.
Furthermore, mid-layer weight changes were found to correlate most strongly with performance improvements, indicating that critical adaptations often take place in these layers. By embedding every model checkpoint into a common latent space, we found that (i) model architecture exerts a stronger influence than the SFT corpus, (ii) training epochs drive diverse runs toward a shared instruction-compatible region, (iii) large instruction sets tend to relocate models toward the periphery—often reducing accuracy relative to smaller sets—and (iv) LoRA trajectories almost coincide with full-parameter ones, diverging only slightly on the periphery; this mirrors the small but systematic trade-off we observed between knowledge-heavy tasks (full-parameter advantaged) and open-ended QA (LoRA advantaged).
@@
Discovering the importance of mid-layer changes could reshape fine-tuning strategies. Updating the mid-layers, or monitoring them closely, could provide more efficient or interpretable SFT. Observing common mid-layer change patterns across models suggests a shared mechanism for task-related knowledge acquisition.
@@
In this study, we conducted comprehensive fine-tuning experiments on pre-trained models in the 7-9B parameter class. Due to limited computational resources, we were unable to extend our verification to larger models (e.g., 70B, 175B, or MoE architectures). It is unclear whether the findings obtained in this research can be generalized to these larger models.
We emphasize the transparency and reproducibility of our analysis, and therefore, we trained our models using only open-access training datasets. Compared to existing publicly available English corpora, there are still limited multilingual instruction-tuning datasets. Consequently, our study used only English training datasets. A comprehensive investigation into cross-lingual knowledge transfer and its differing effects remains a subject for future work. We use about 10 popular training datasets for SFT, possibly limiting generality for highly specialized tasks or broader multilingual corpora.
While perplexity proved insightful, it can fluctuate based on tokenizer design and base training distributions, indicating a need for more nuanced measures.
@@
We will release over 1,000 fine-tuned models as part of this study. While we do not anticipate major risks, we acknowledge the potential for misuse—such as generating harmful or misleading content. To mitigate this, all released models will include a responsible use clause and detailed model cards describing limitations. We encourage responsible use.
@@ Appendix B Description of Training Datasets
Classic
NLP
FLAN Knowledge Open-Orca/FLAN 226,575 21.69
@@
FLAN Collection (Wei et al., 2022) is the remix file flan2021_zsnoopt_submix_data.json. Specifically, it corresponds to the FLAN‑2021 sub‑mix and employs the zero‑shot, no‑options template variant (i.e., prompts contain only the instruction without in‑context examples or candidate options). We follow the taxonomy of Dubey et al. (2024); Contributors (2023) and split the data into three thematic subsets.
@@ Appendix C Preliminary Experiments
This combination of hyperparameters resulted in 96 unique experimental conditions. For each condition, we trained for 10 epochs and saved a checkpoint after every epoch (epochs 1 to 10). We treated these per-epoch checkpoints as distinct candidate models, yielding 960 models per dataset. Given that we utilized six datasets, the total number of trained models reached 5,760, consuming at least 23,040 GPU-hours.
@@ Appendix D Description of Training Settings
The computational time for fine-tuning on 1k samples varied depending on the model, batch size, and training method, but on average, each run took approximately 30 minutes. To accelerate training, we employed Flash Attention 2 (Dao, 2023) and DeepSpeed (Rasley et al., 2020) for all models. Training these 1,059 runs required at least 530 GPU-hours.